# Reweight-annealing method for calculating the value of partition function via quantum Monte Carlo

### v3

@@ Abstract
Efficient and accurate algorithm for partition function, free energy and thermal entropy calculations is of great significance in statistical physics and quantum many-body physics. Here we present an unbiased but low-technical-barrier algorithm within the quantum Monte Carlo framework, which has exceptionally high accuracy and no systemic error. Compared with the conventional specific heat integral method and Wang-Landau sampling algorithm , our method can obtain a much more accurate result of the sub-leading coefficient of the entropy consistent with CFT. This method can be widely used in both classical and quantum Monte Carlo simulations and is easy to be parallelized on computer.
@@
On the other hand, the high-precision PF and its related quantities, such as free energy and entropy, are significantly important in numerical calculations since it means more effective information. For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain PF would be crucial, because its (high order) derivative reveals potential phase transitions. The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice Snyder et al. (2001); Skjærvø et al. (2020); Udagawa et al. (2021); Bramwell and Harris (2020); Udagawa et al. (2021), topological order Jiang et al. (2012); Castelnovo et al. (2012); Sachdev (2018); Levin and Wen (2006); Yan et al. (2021, 2022), fracton phases Pretko et al. (2020); Nandkishore and Hermele (2019); Gromov and Radzihovsky (2024); Zhou et al. (2022), etc. Besides, the precision of entropy can even reveal important intrinsic physics. In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al. (2020); Wu et al. (2021a); Zhao et al. (2021); Wu et al. (2021b). Even a small numerical error may make our concerned quantities deviate a lot because the sub-leading term is much more sensitive than the leading one.
@@
In this paper, we present a more accurate and efficient method for calculating PF, dubbed reweight-annealing algorithm, within the framework of Monte Carlo (MC) simulation, either quantum or classical. It is inspired by thermal and quantum annealing algorithm in the quantum simulation area Huse and Fisher (1986); Santoro et al. (2002); Lucas (2014); Kadowaki and Nishimori (1998); Santoro et al. (2002); Mézard and Montanari (2009); Heim et al. (2015); Yan et al. (2023); Ding et al. (2023) and recent developments of the incremental trick in the entanglement calculations of QMC Alba (2017); D’Emidio (2020); Zhao et al. (2022a, b); Pan et al. (2023); Zhou et al. (2024). Our method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer.
@@
Basic idea.- Given a partition function $Z({\mathbf{p}})$, where ${\mathbf{p}}=[p_{1},p_{2},\cdots]$ denotes the parameters in the PF such as the temperature $T$ and some Hamiltonian parameter $s$. For example, Fig.1 (a) shows a diagram of the two parameters. For two different parameter points ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, we define the reweighting ratio which can be measured in Monte Carlo
$\frac{Z({\mathbf{p}}^{\prime})}{Z({\mathbf{p}}^{\prime\prime})}=\bigg{\langle}\frac{W({\mathbf{p}}^{\prime})}{W({\mathbf{p}}^{\prime\prime})}\bigg{\rangle}_{{\mathbf{p}}^{\prime\prime}}$ (1)
@@
where $W({\mathbf{p}}^{\prime})$ and $W({\mathbf{p}}^{\prime\prime})$ denote the weight of the same configuration under different parameters ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, and the subscript indicates the result comes from the MC simulation under parameter ${\mathbf{p}}^{\prime\prime}$. As an example, we take ${\mathbf{p}}=\beta$, where $\beta$ denotes the inverse temperature. Under the framework of stochastic series expansion (SSE), we can then derive $Z(\beta^{\prime})/Z(\beta^{\prime\prime})=\langle(\beta^{\prime}/\beta^{\prime\prime})^{n}\rangle_{\beta^{\prime\prime}}$, where $n$ is the number of non-identity operators in the SSE simulation. The details of the $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ have been explained in the supplemental materials (SM).
@@
The second step is to identify a reference point ${\mathbf{p}}_{0}$, of which the value of $Z({\mathbf{p}}_{0})$ has already been known. Therefore, by evaluating $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$, we can directly determine the value of $Z({\mathbf{p}}^{\prime\prime})$. In the example that ${\mathbf{p}}=\beta$, apparently $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the degree of freedom of a single particle respectively.
However, when ${\mathbf{p}}_{0}$ and ${\mathbf{p}}^{\prime\prime}$ are far away in the parameter space, the value of $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$ would be close to zero or infinity, resulting in the problem of inefficiency and low precision for the MC simulation. This can be easily understood in the reweighting frame as shown in Fig. 1 (b). If we want to use a well-known distribution $Z({\mathbf{p}}^{\prime\prime})=\sum W({\mathbf{p}}^{\prime\prime})$ to calculate another distribution $Z({\mathbf{p}}^{\prime})=\sum W({\mathbf{p}}^{\prime})$ via resetting the weight of samplings, the weight before/after the resetting of the same configuration should be close to each other. It is still an importance sampling in this sense and requires that ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$ should be close enough.
@@
$\frac{Z({\mathbf{p}}_{0})}{Z({\mathbf{p}}^{\prime\prime})}=\prod_{k=1}^{m}\frac{Z({\mathbf{p}}_{k-1})}{Z({\mathbf{p}}_{k})}$ (2)
@@
where ${\mathbf{p}}_{m}\equiv{\mathbf{p}}^{\prime\prime}$ and $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}_{k}})\to 1$ as long as ${\mathbf{p}}_{k}\to{\mathbf{p}}_{k-1}$. Here, ${\mathbf{p}}$ can be perceived to gradually change from the point of interest ${\mathbf{p}}^{\prime\prime}$ to the reference point ${\mathbf{p}}_{0}$ along a trajectory (annealing path) in the parameter space, akin to an adiabatic process or annealing as shown in Fig.1 (a). Therefore by calculating Eq. (2), not only we are able to achieve a higher precision of $Z({\mathbf{p}}^{\prime\prime})$, but concurrently track the variation of PF along this annealing path. This allows us to further analyze the variation of the free energy (and its derivatives) of the system.
Apparently, the annealing path that ${\mathbf{p}}$ follows determines how large $m$ should we need and how ${\mathbf{p}}$ varies to perform the calculations of Eq. (2). It can be expected that, similar to conventional thermal annealing or quantum annealing, should the annealing trajectory approach a critical point, an increased number of divisions would be necessary in its vicinity. However, we point out that this actually affects little to our algorithm since numerical simulations only deal with finite system, where the criticality does not result in significant inefficiency. We have also verified this point in our simulations. For example, in the 2D spin-$1/2$ XXZ model which we present below, where ${\mathbf{p}}=T$ is the temperature, we implemented an equal division strategy with step size $\Delta T=0.01$. With $20000$ Monte Carlo samples for each bin, we are able to achieve a sufficiently smooth curve for further numerical differentiation. More discussions about the segmentation are in the section of “Dynamical annealing process”.
@@
Thermal and quantum reweight-annealing- As it shown in Fig. 1, if ${\mathbf{p}}=\beta$, as previously mentioned, the annealing path connects an unknown inverse temperature point $\beta^{\prime\prime}$ to the infinite temperature point $\beta_{0}$, which serves as the reference. For convenience, we call this kind of process the thermal reweight-annealing (T-Re-An). Similarly, we can consider the so-called quantum reweight-annealing (Q-Re-An) by setting ${\mathbf{p}}=s$, where $s$ comes from $H(s)=sH_{0}+(1-s)H_{1}$, which is the archetypal Hamiltonian of quantum annealing with $Z(s)=e^{-\beta H(s)}$. Here, $H_{1}$ correspond to our reference point $s=0$, and we would like to obtain the PF related to $H_{0}$. We emphasis that different from the thermal version, the choice of $H_{1}$ here is much more flexible. For example, we can take $H_{1}=\sum_{i}\sigma_{i}^{x}$. Then at the zero temperature ($\beta\to\infty$), $H_{1}$ only has a unique ground state, thus its thermal entropy is minimized to be zero, resulting in $\ln Z(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$. More details about the T-Re-An and Q-Re-An scheme can be found in the SM.
@@
Dynamical annealing process.- For the thermal and quantum reweight-annealing schemes, we provide a dynamical strategy to automatically regulate the alteration of ${\mathbf{p}}$ throughout the annealing process. This not only serves as an economic way for the division in Eq. (2), but can avoid the problem of precision loss when different $r({\mathbf{p}}_{k-1},{\mathbf{p}}_{k})$ differ by several orders of magnitude in the computer. For demonstration, we only discuss the thermal case, i.e. ${\mathbf{p}}=\beta$, in this section.
Without loss of generality, we assume $\beta_{k-1}<\beta_{k}$ in Eq. (2), and define the rate $\alpha_{k}\equiv\beta_{k-1}/\beta_{k}<1$. Within the framework of SSE, we have
@@
$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\langle\alpha^{n_{k}}_{k}\rangle_{\beta_{k}}$ (3)
where $n_{k}$ is the number of all non-identity operators in the SSE simulation, and the subscript $\beta_{k}$ denotes that this simulation is performed at the parameter point $\beta_{k}$.
The annealing process starts from $k=m$, and we require $Z(\beta_{k-1})/Z(\beta_{k})=\epsilon<1$, where the choice of $\epsilon$ should match $N_{\text{MC}}$, the number of MC steps we use for samplings, as the error in MC simulations is proportional to $1/\sqrt{N_{\text{MC}}}$. Next, we set $n_{k}\approx\tilde{n}_{k}=\Lambda_{k}\beta_{k}L$ ($\Lambda_{k}\beta_{k}L^{2}$) for 1D (2D) system, and $\Lambda_{k}$ can be ascertained through the examination of result convergence. This comes from the fact that in SSE, the energy satisfies $E_{k}=-\langle n_{k}\rangle/\beta_{k}$. Practically, we set $\Lambda_{k}$ to be the same for all $k$ for convenience. This strategy not only makes $Z(\beta_{k-1})/Z(\beta_{k})\approx\epsilon$ for all $k$, but determines the next step in the annealing path, i.e. $\beta_{k-1}=\alpha_{k}\beta_{k}$. The division will be more sparse when $\beta$ gets smaller as we expected, where the thermal fluctuation has destroyed any order and $Z(\beta_{k-1})/Z(\beta_{k})\to 1$, which has little contribution to the result. Apparently, this strategy naturally allows for parallelization of evaluating different $Z(\beta_{k-1})/Z(\beta_{k})$ in Eq. (2). It is also easy to calculate that the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$ for large $L$ or $(\beta-\beta_{0})$, where $\Lambda_{k}=\Lambda$ for all $k$. Therefore, this dynamical annealing strategy does not require exponentially increasing number of divisions.

### v5

# Reweight-annealing method for evaluating the partition function via quantum Monte Carlo calculations

@@ Abstract
Efficient and accurate algorithm for partition function, free energy and thermal entropy calculations is of great significance in statistical physics and quantum many-body physics. Here we present an unbiased but low-technical-barrier algorithm within the quantum Monte Carlo framework, which has exceptionally high accuracy and no systemic error. Compared with the conventional specific heat integral method and Wang-Landau sampling algorithm, our method can obtain a much more accurate result of the sub-leading coefficient of the entropy consistent with theory. This method can be widely used in both classical and quantum Monte Carlo simulations and is easy to be parallelized on computer.
I Introduction
@@
On the other hand, the high-precision PF and its related quantities, such as free energy and entropy, are significantly important in numerical calculations since it means more effective information. For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain PF would be crucial, because its (high order) derivative reveals potential phase transitions. The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice [20, 21, 22, 23, 22, 24], topological order [25, 26, 27, 28, 29, 30], fracton phases [31, 32, 33, 34], etc. Besides, the precision of entropy can even reveal important intrinsic physics. In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. Even a small numerical error may make our concerned quantities deviate a lot because the sub-leading term is much more sensitive than the leading one.
@@
In this paper, we present a more accurate and efficient method for calculating PF, dubbed reweight-annealing algorithm, within the framework of Monte Carlo (MC) simulation, either quantum or classical. It is inspired by thermal and quantum annealing algorithm in the quantum simulation area [46, 47, 48, 49, 47, 50, 51, 52, 53, 54] and recent developments of the incremental trick in the entanglement calculations of QMC [55, 56, 57, 58, 59, 60], while the earliest idea can date back to the thermodynamic integration [61, 62, 63], and many of its variants found the applications in both condensed matter physics and high energy physics [64, 65, 66, 67, 68, 69, 70]. The method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer. With the annealing scheme we introduce below, a polynomial complexity can be proved, enabling better control over the computational precision for large systems.
This paper is arranged as follows. In Sec. II, we elaborate the algorithmic and technical details of our method, including the annealing scheme, then to demonstrate the efficacy of the method, we provide three examples in Sec. III. In the end, a summary and discussions are put in Sec. IV.
II Method
II.1 Basic idea
@@
Given a partition function $Z(\vec{p})$, where $\vec{p}=[p_{1},p_{2},\cdots]$ denotes the parameters in the PF such as the temperature $T$ and some Hamiltonian parameter $s$. For example, Fig.1 (a) shows a diagram of the two parameters. For two different parameter points $\vec{p}^{\prime}$ and $\vec{p}^{\prime\prime}$, we define the reweighting ratio which can be measured in Monte Carlo
$\frac{Z(\vec{p}^{\prime})}{Z(\vec{p}^{\prime\prime})}=\bigg{\langle}\frac{W(\vec{p}^{\prime})}{W(\vec{p}^{\prime\prime})}\bigg{\rangle}_{\vec{p}^{\prime\prime}}$ (1)
@@
where $W(\vec{p}^{\prime})$ and $W(\vec{p}^{\prime\prime})$ denote the weight of the same configuration under different parameters $\vec{p}^{\prime}$ and $\vec{p}^{\prime\prime}$, and the subscript indicates the result comes from the MC simulation under parameter $\vec{p}^{\prime\prime}$. As an example, we take $\vec{p}=\beta$, where $\beta$ denotes the inverse temperature. Under the framework of stochastic series expansion (SSE)[1, 2], we can then derive $Z(\beta^{\prime})/Z(\beta^{\prime\prime})=\langle(\beta^{\prime}/\beta^{\prime\prime})^{n}\rangle_{\beta^{\prime\prime}}$, where $n$ is the number of non-identity operators in the SSE simulation. The details of the $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ can be found in Appx. A.
@@
The second step is to identify a reference point $\vec{p}_{0}$, of which the value of $Z(\vec{p}_{0})$ has already been known. Therefore, by evaluating $Z(\vec{p}_{0})/Z(\vec{p}^{\prime\prime})$, we can directly determine the value of $Z(\vec{p}^{\prime\prime})$. In the example that $\vec{p}=\beta$, apparently $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the degree of freedom of a single particle respectively.
However, when $\vec{p}_{0}$ and $\vec{p}^{\prime\prime}$ are far away in the parameter space, the value of $Z(\vec{p}_{0})/Z(\vec{p}^{\prime\prime})$ would be close to zero or infinity, resulting in the problem of inefficiency and low precision for the MC simulation. This can be easily understood in the reweighting frame as shown in Fig. 1 (b). If we want to use a well-known distribution $Z(\vec{p}^{\prime\prime})=\sum W(\vec{p}^{\prime\prime})$ to calculate another distribution $Z(\vec{p}^{\prime})=\sum W(\vec{p}^{\prime})$ via resetting the weight of samplings, the weight before/after the resetting of the same configuration should be close to each other. It is still an importance sampling in this sense and requires that $\vec{p}^{\prime}$ and $\vec{p}^{\prime\prime}$ should be close enough [62, 71].
@@
$\frac{Z(\vec{p}_{0})}{Z(\vec{p}^{\prime\prime})}=\prod_{k=1}^{m}\frac{Z(\vec{p}_{k-1})}{Z(\vec{p}_{k})}$ (2)
@@
where $\vec{p}_{m}\equiv\vec{p}^{\prime\prime}$ and $Z(\vec{p}_{k-1})/Z(\vec{p_{k}})\to 1$ as long as $\vec{p}_{k}\to\vec{p}_{k-1}$. Here, $\vec{p}$ can be perceived to gradually change from the point of interest $\vec{p}^{\prime\prime}$ to the reference point $\vec{p}_{0}$ along a trajectory (annealing path) in the parameter space, akin to an adiabatic process or annealing as shown in Fig.1 (a). Therefore by calculating Eq. (2), not only we are able to achieve a higher precision of $Z(\vec{p}^{\prime\prime})$, but concurrently track the variation of PF along this annealing path. This allows us to further analyze the variation of the free energy (and its derivatives) of the system.
II.2 Thermal and quantum reweight-annealing
@@
As it shown in Fig. 1, if $\vec{p}=\beta$, as previously mentioned, the annealing path connects an unknown inverse temperature point $\beta^{\prime\prime}$ to the infinite temperature point $\beta_{0}$, which serves as the reference. For convenience, we call this kind of process the thermal reweight-annealing (T-Re-An). Similarly, we can consider the so-called quantum reweight-annealing (Q-Re-An) by setting $\vec{p}=s$, where $s$ comes from $H(s)=sH_{0}+(1-s)H_{1}$, which is the archetypal Hamiltonian of quantum annealing with $Z(s)=e^{-\beta H(s)}$. Here, $H_{1}$ correspond to our reference point $s=0$, and we would like to obtain the PF related to $H_{0}$. Different from the thermal version, the choice of $H_{1}$ here is much more flexible. For instance, we can take $H_{1}=\sum_{i}\sigma_{i}^{x}$. Then at the zero temperature ($\beta\to\infty$), $H_{1}$ only has a unique ground state, thus its thermal entropy is minimized to be zero, resulting in $\ln Z(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$. More details about the T-Re-An and Q-Re-An scheme can be found in the SM.
II.3 Automatic annealing process
@@
Here we provide a scheme to automatically regulate the alteration of $\vec{p}$ throughout the annealing process. This not only serves as an economic way for the division in Eq. (2), but can avoid the problem of precision loss when different $Z(\vec{p}_{k-1})/Z(\vec{p}_{k})$ differ by several orders of magnitude in the computer. For demonstration, we discuss the thermal case, i.e. $\vec{p}=\beta$ in this section, and the case that $\vec{p}=s$ is similar.
For $\beta_{0}<\beta$, when $\beta_{0}$ and $\beta$ are far away in the parameter space, to ensure importance sampling, we divide the interval $[\beta_{0},\beta]$ into $m$ subintervals $[\beta_{0},\beta_{1}]$, $[\beta_{1},\beta_{2}]$, $\cdots$, $[\beta_{m-1},\beta_{m}]$, where $\beta_{m}\equiv\beta$, i.e.
as Eq. (2). Each subinterval $[\beta_{k-1},\beta_{k}]$ ought to be sufficiently narrow so that the distribution at $\beta_{k}$ can well samples the configurations at $\beta_{k-1}$. Otherwise, the sampling process would be inefficient. Then the natural questions following are how to partition $[\beta_{0},\beta]$ and the how many subintervals we need for Eq. (3).
One crude way is just dividing $[\beta_{0},\beta]$ into equal subintervals or in a geometric progression. As long as the division number is large enough, the result would be accurate. However, this way is not economic. We can expect that when $\beta\to 0$ (approaching infinite high temperature), the division can be more sparse compared to that at low temperatures. Another worry is that if $\beta_{k}$ and $\beta_{k-1}$ are separated on the two sides of the critical point, we may need a bulk of samples to estimate $Z(\beta_{k-1})/Z(\beta_{k})$.
In addition to the division problem, $Z(\beta_{0})/Z(\beta)$ could be an exponentially small number exceeding the data range of computer (e.g. $1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data), and we should store $\ln[Z(\beta_{0})/Z(\beta)]$ in the computer. However, simply applying the logarithm to Eq. (3) and converting the multiplication to addition may lead to precision loss in computer calculations if $\ln[Z(\beta_{k-1})/Z(\beta_{k})]$ and $\ln[Z(\beta_{k})/Z(\beta_{k+1})]$ differ by several orders of magnitude.
To address the two problems above, we present the automatic/pseudo-automatic annealing scheme. The automatic annealing scheme starts by fixing $Z(\beta_{k-1})/Z(\beta_{k})=\epsilon<1$ for all $k$. We will talk about the choice of $\epsilon$ later, and first assume we have already selected some $\epsilon$.
Recall that in SSE, the energy estimator is $E(\beta_{k})=-{\langle n\rangle_{\beta_{k}}}/{\beta_{k}}$, where $n_{k}\equiv\langle n\rangle_{\beta_{k}}$ denotes the average number of non-identity operators at temperature $\beta_{k}$, which itself is also an observable that can be estimated in SSE simulations. After we have achieved an estimation $\tilde{n}_{k}$ of $n_{k}$ from simulations (this is exactly same with measuring energy by counting the number of non-identity operators in each SSE configuration), we can bring the result into $Z(\beta_{k-1})/Z(\beta_{k})$ to have
@@
$\epsilon=\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n}\bigg{\rangle}_{\beta_{k}}\approx\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{\tilde{n}_{k}}$ (4)
Attention that Eq. (4) is just an approximation, but it gives us a way to divide $[\beta_{0},\beta]$.
Imagine that we start from the $\beta\equiv\beta_{m}$, which is the right end point of $[\beta_{0},\beta]$, then using Eq. (4), we can determine the position (value) of $\beta_{m-1}$, which is
$\beta_{m-1}=\epsilon^{1/\tilde{n}_{k}}\beta_{m}$ (5)
Similarly, we can obtain $\beta_{m-2},\beta_{m-3},\cdots$, and it stops when we encounter some $\beta_{j}$ such that $\beta_{j-1}<\beta_{0}$. Then $\beta_{j}$ is set to be $\beta_{1}$ and the division procedure is over. Notice that each $\beta_{k-1}$ must rely on the knowledge of $\beta_{k}$, and each $\tilde{n}_{k}$ is obtained from the corresponding SSE simulations. This division procedure is similar to a standard (thermal or quantum) annealing process, in which we start from an initial parameter point, and step-by-step change it to a target parameter point. Here we go from $\beta\equiv\beta_{m}$ to $\beta_{0}$. Of course one can inversely go from $\beta_{0}$ to $\beta$ by considering $Z(\beta)/Z(\beta_{0})$ rather than $Z(\beta_{0})/Z(\beta)$.
About the choice of $\epsilon$, since we have to estimate $Z(\beta_{k-1})/Z(\beta_{k})$, we hope this value smaller than $1$ but close to $1$ so that the distribution at $\beta_{k-1}$ and $\beta_{k}$ would have big overlap to make the simulations more efficient. Just as we perform some common MC simulations to estimate some observable, the MC steps and the number of bins required for an accurate estimation should be tried. Here in our algorithm, it is similar actually. After we fix some $\epsilon$, we have to determine the required MC steps and bins via some trial simulations. We can expect that if we take a very small $\epsilon$, the MC steps required for an accurate estimation would be large. However, no matter what $\epsilon$ we choose, in principle, as long as we have enough MC steps and bins, the estimation of $Z(\beta_{k-1})/Z(\beta_{k})$ can be accurate. If $\epsilon$ is very close to $1$, we would also have some problems because the number of divided segments $m$ would be very large. Therefore there is some tradeoff about the choice of $\epsilon$.
II.4 Pseudo-automatic annealing process
The “automatic annealing” scheme mentioned above can not be parallelized since we have to achieve each $\tilde{n}_{k}$ every time from the simulations so that we can go to the next point. Therefore, practically, we give a priori estimate of the value of $\tilde{n}_{k}$ without doing the simulations. This is what we call the pseudo-automatic annealing scheme used in our paper.
Recall that energy is an extensive quantity, then it should satisfy $E\sim L^{d}$ for a $d$-dimensional system. In SSE, we then have $\langle n\rangle_{\beta_{k}}\sim\beta_{k}L^{d}$. Here $\beta_{k}$, $L$ and $d$ are given, therefore to obtain $\langle n\rangle_{\beta_{k}}$, we need to know some prefactor $\Lambda_{k}$ such that $\langle n\rangle_{\beta_{k}}\simeq\Lambda_{k}\beta_{k}L^{d}\approx\tilde{n}_{k}$. It might be questioned that not even we have to decide the value of $\epsilon$, but we have another hyperparameter $\Lambda_{k}$ to determine. In fact, we still have one since now Eq. (4) becomes
$\epsilon\approx\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{\Lambda_{k}\beta_{k}L^{d}}$ (6)
If we substitute the real $\Lambda_{k}$ with some incorrect $\tilde{\Lambda}_{k}$ we manually set, this just effectively changes the value of $Z(\beta_{k-1})/Z(\beta_{k})$ from $\epsilon$ to some $\epsilon_{\text{eff}}$. Therefore we can fix $\Lambda_{k}$ to be some value such as $L$ in our simulations for all $k$ for convenience, then the problem reduces to choosing a good $\epsilon$. The uniform choice of $\Lambda_{k}\equiv\Lambda$ would make different $Z(\beta_{k-1})/Z(\beta_{k})$ slightly deviated from our desired ratio $\epsilon_{\text{eff}}$, but it is okay since we always have to determine the MC steps and bins for estimation. In this aspect, there is no different no matter for $\epsilon_{\text{eff}}$ (with small deviations at each point) or the $\epsilon$ in the “automatic annealing” scheme in the last subsection.
Next we show that this pseudo-automatic annealing scheme actually requires the number of divisions scales proportional to the system size and the inverse temperature for large systems. For the 1D system as an example,
or
$\beta_{k}-\beta_{k-1}=-\frac{\ln\epsilon}{\Lambda L}$ (8)
then the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$. Notice that we have ignored the case that $\beta_{k}\to 0$ in Eq.(7). This is because when $\beta\to 0$, we have $\epsilon^{1/(\Lambda\beta_{k}L)}\to 0$ ($\epsilon<1$), then the division becomes sparser and has little contributions to the total number of divisions. Another thing worthy of attention is that when $\beta\to 0$, the small value of $\epsilon^{1/(\Lambda\beta_{k}L)}\to 0$ would also exceed the range of a number saved in the computer. Therefore we manually set a lower bound for $\alpha_{k}$ (e.g. $\min{\alpha_{k}}=\epsilon$ in our simulation). The existence of this lower bound makes the division in the region that $\beta\to 0$) slightly denser than that Eq. (6) requires. Since most $Z(\beta_{k-1})/Z(\beta_{k})$ still at the same order of magnitude, and the more divisions when $\beta\to 0$ also only constitutes a very small proportion. This affects little to the precision as well as the number of divisions calculated in Eq. (8). In Fig. 2, we plot the number of divisions in the T-Re-An method of first example (see Sec. III.1) in our paper for reference, where we take $\Lambda=L$, and compare it with the estimation number $(\beta-\beta_{0})\Lambda L/|\ln(\epsilon)|$.
Figure 2: The number of divisions in the T-Re-An method of first example (see Sec. III.1), where we take $\epsilon=0.01$, $\beta=30$ and $\Lambda=L$.
III Examples
III.1 1D spin-1/2 Heisenberg chain
As the first example, we consider an 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) with $\beta=30$ and length $L$. We use both the T-Re-An method and the Q-Re-An method to calculate the PF. For the quantum version, we write
where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}\vec{S}_{i}\cdot\vec{S}_{i+1}$. Apparently, the unique ground state of $H_{1}$ is a tensor product of $L/2$ singlets, each carrying energy $-3/4$ ($\hbar=1$). The estimator for $Z(s^{\prime})/Z(s^{\prime\prime})$ is $\langle(s^{\prime}/s^{\prime\prime})^{n_{\text{odd}}}\rangle_{s^{\prime\prime}}$ in this model (see the derivation in Appx. A), with $n_{\text{odd}}$ denoting the amount of non-identity operators acting on all the bonds with odd indices within the framework of SSE.